OpenStudy (anonymous):

if (2x+1)( x-5) =2(x^2-1), what is the value of x?

OpenStudy (anonymous):

wouldn't 2(x^2-1) be distributed to 2x^2-2?

OpenStudy (anonymous):

oops oh sorry

OpenStudy (anonymous):

x = -3/9=-1/3

OpenStudy (anonymous):

do you need work?????????

OpenStudy (anonymous):

yes i didn't know to cancel the 2x^2 so your first answer got me on the right track

OpenStudy (anonymous):

\[2x^2-9x-5=2x^2-2\] \[-5+2=2x^2-2x^2+9x\] \[9x=-3\] \[x=-3/9\] \[x=-1/3\]

OpenStudy (anonymous):

oh! i did not know what i meant by work thanks anyways